# Adaptive Collaborative Correlation Learning-based Semi-Supervised Multi-Label Feature Selection

### v3

@@ 4 Optimization and Algorithm
$\min_{\mathbf{W},\mathbf{F},\mathbf{S},\mathbf{P}}\|\mathbf{H}\mathbf{X}^{T}\mathbf{W}-\mathbf{H}\mathbf{F}\|_{F}^{2}+{\|{\mathbf{F}_{l}-\mathbf{Y}_{l}}\|_{F}^{2}}+\lambda\|\mathbf{W}\|_{2,1}$ (8)
$~{}~{}\quad\quad\quad+\theta(\operatorname{Tr}(\mathbf{F}^{T}\mathbf{L}_{s}\mathbf{F})+\operatorname{Tr}(\mathbf{W}^{T}\mathbf{X}\mathbf{L}_{s}\mathbf{X}^{T}\mathbf{W})+\alpha\|\mathbf{S}\|_{F}^{2})$
$~{}~{}\quad\quad\quad+\!\mu(\operatorname{Tr}(\mathbf{F}\mathbf{L}_{p}\mathbf{F}^{T})+\beta\|\mathbf{P}\|_{F}^{2})$
@@
$~{}~{}~{}~{}~{}~{}~{}p_{ij}\geq 0,\mathbf{P}\mathbf{1}_{c}=\mathbf{1}_{c}.$
@@ 4.1 Update $\mathbf{W}$ by Fixing Other Variables
4.2 Update $\mathbf{F}$ by Fixing Other Variables
$\min_{\mathbf{F}}\|\mathbf{H}\mathbf{X}^{T}\mathbf{W}-\mathbf{H}\mathbf{F}\|_{F}^{2}+\|{\mathbf{F}_{l}-\mathbf{Y}_{l}}\|_{F}^{2}+\theta\operatorname{Tr}(\mathbf{F}^{T}\mathbf{L}_{s}\mathbf{F})$ (15)
$\quad\quad+\mu\operatorname{Tr}(\mathbf{F}\mathbf{L}_{p}\mathbf{F}^{T}).$
$\mathbf{Q}\mathbf{F}+\mu\mathbf{F}\mathbf{L}_{p}=\mathbf{C}$ (17)
Input: Data matrix $\mathbf{X}$, centering matrix $\mathbf{H}$, predicted label matrix $\mathbf{F}$, Laplacian matrix $\mathbf{L}_{s}$, and the parameters $\lambda$ and $\theta$.
@@ 4.3 Update $\mathbf{S}$ by Fixing Other Variables
While $\mathbf{W}$, $\mathbf{F}$ and $\mathbf{P}$ are fixed, the optimization problem in Eq. (8) is equivalent to solve Eq. (LABEL:New3). By setting a matrix $\mathbf{M}=(m_{ij})_{n\times n}$, where $m_{ij}=\frac{1}{2}\|\mathbf{f}_{i.}-\mathbf{f}_{j.}\|_{2}^{2}+\frac{1}{2}\|\mathbf{W}^{T}\mathbf{x}_{.i}-\mathbf{W}^{T}\mathbf{x}_{.j}\|_{2}^{2}$, Eq. (LABEL:New3) is transformed into
@@
In practice, each sample tends to be similar to its neighbours. Therefore, we aim for $\mathbf{s}_{i.}$ to have $k_{s}$ nonzero values, indicating that the $i$th sample is similar to its $k_{s}$ closest neighbours. Assuming that $m_{i1},m_{i2},\cdots,m_{in}$ are arranged in ascending order, we prefer ${s}_{ik_{s}}>0$ and ${s}_{ik_{s}+1}=0$. As a result, we have $\alpha=\frac{m_{ik_{s}+1}}{2\psi}$. Besides, according to the constraint $\mathbf{s}_{i.}\mathbf{1}_{n}=1$, we get $\psi=\frac{1}{k_{s}}+\frac{\sum_{j=1}^{k_{s}}m_{ij}}{2k_{s}\alpha}$. By substituting $\psi$ into the obtained $\alpha$, we can calculate the regularization parameter as $\alpha=\frac{k_{s}m_{ik_{s}+1}-\sum_{j=1}^{k_{s}}m_{ij}}{2}$. Then, we achieve the optimal solution ${s}_{ij}$ as follows:
${s}_{ij}=\left\{\begin{array}[]{cl}\frac{m_{ik_{s}+1}+m_{ij}}{k_{s}m_{ik_{s}+1}-\sum_{h=1}^{k_{s}}m_{ih}}&j\leq k_{s};\\ 0&j>k_{s}.\end{array}\right.$ (21)
@@ 4.4 Update $\mathbf{P}$ by Fixing Other Variables
$\min_{\mathbf{P}}\frac{1}{2}\sum_{i,j=1}^{c}\|\mathbf{f}_{.i}-\mathbf{f}_{.j}\|_{2}^{2}p_{ij}+\beta\sum_{i=1}^{c}\|\mathbf{p}_{i.}\|_{2}^{2}$ (22)
@@
Following a similar procedure to that used in solving Eq. (19), we can determine the regularization parameter $\beta=\frac{k_{p}g_{ik_{p}+1}-\sum_{j=1}^{k_{p}}g_{ij}}{2}$, where $k_{p}$ denotes the number of nonzero elements in $\mathbf{p}_{i.}$. Furthermore, the optimal solution ${p}_{ij}$ is formulated as follows:
$p_{ij}=\left\{\begin{array}[]{cl}\frac{g_{ik_{p}+1}+g_{ij}}{k_{p}g_{ik+1}-\sum_{h=1}^{k_{p}}g_{ih}}&j\leq k_{p};\\ 0&j>k_{p}.\end{array}\right.$ (25)
@@ 5.1 Convergence Analysis
Due to the objective function in Eq. (8) not being convex w.r.t. four variables $\mathbf{W}$, $\mathbf{F}$, $\mathbf{S}$ and $\mathbf{P}$ simultaneously, we address this by dividing it into four sub-objective functions, namely, Eqs. (9), (15), (18), and (22), each of which is separately solved with other variables fixed. Therefore, the convergence of Algorithm 2 can be demonstrated by proving that each sub-objective function decreases monotonically. We first present the following theorem to prove the monotonic decrease of the objective function in Eq. (9) when updating $\mathbf{W}$ with $\mathbf{F}$, $\mathbf{S}$ and $\mathbf{P}$ are fixed.
@@
We can prove this theorem based on the property in [33] and the result in [34]. Due to space constraints, the detailed proof of Theorem 1 is provided in the supplementary material.
Besides, as Eq. (15) is equivalent to Eq. (16), we only need to demonstrate that the objective function in Eq.(16) with respect to ${\mathbf{F}}$ is a monotonically decreasing function. It is easy to verify that the matrix $\mathbf{Q}$ and the Laplacian matrix $\mathbf{L}_{p}$ are both positive semidefinite. Consequently, the quadratic forms $\mathbf{F}^{T}\mathbf{Q}\mathbf{F}-2\mathbf{F}^{T}\mathbf{C}$ and $\mathbf{F}\mathbf{L}_{p}\mathbf{F}^{T}$ are convex. Hence, the objective function in Eq.(16) is convex. According to [35], we can easily obtain that the objective function in Eq.(16) monotonically decrease by using the updating rules in Algorithm 2. Additionally, the objective functions in Eqs. (18) and (22) are convex w.r.t. $\mathbf{S}$ and $\mathbf{P}$, and possess closed-form solutions, ensuring the convergence of the updates for $\mathbf{S}$ and $\mathbf{P}$, respectively. Therefore, we can conclude that the objective function in Eq.(8) monotonically decreases with each iteration of Algorithm 2. Moreover, the convergence behavior of Algorithm 2 is further demonstrated through experiments.
@@ 5.2 Complexity Analysis
In Algorithm 2, four variables are updated alternately by solving Eqs. (9), (15), (18), and (22). The optimization of Eq. (9) is further transformed into solve Eq. (13). According to [34], the time complexity for updating $\mathbf{W}$ in Eq. (13) is $\mathcal{O}(d^{3})$. For updating $\mathbf{F}$, a proposed algorithm in [35] is applied into solve Eq. (15). Based on [35], we can calculate the computational complexity for updating $\mathbf{F}$, which is $\mathcal{O}(n^{2}c+c^{2}n)$. Eq. (18) is solved by updating $\mathbf{S}$ row by row. According to the closed form solution of $\mathbf{S}$ in Eq. (21), the time complexity of updating $\mathbf{S}$ is $\mathcal{O}(nk_{s}d)$. The optimization $\mathbf{P}$ in Eq. (22) is similar to solving $\mathbf{S}$. In the same way, we can determine the time complexity of $\mathbf{P}$ as $\mathcal{O}(ck_{p}n)$. Therefore, the total computational complexity of Algorithm 2 is $\mathcal{O}((d^{3}+n^{2}c+c^{2}n+nk_{s}d+ck_{p}n)T)$, where $T$ is the number of iteration.
@@ 6.1.1 Datasets
In our experiments, we collected eight real-world benchmark multi-label datasets, each with distinct statistical characteristics, to present the performance of various feature selection methods. These datasets consist of the biology data set VirusGO  [^1] http://www.uco.es/kdis/mllresources, two medicine datasets Parkinson [^2] https://github.com/XSilverBullet/parkinson- and Medical [^3] http://www.uco.es/kdis/mllresources, as well as five text datasets: Langlog3, Enron3, Slashdot3, Recreation3 and Computer3. The detailed statistics for these multi-label datasets are summarized in Table I. Within this table, Label card. refers to the average number of labels per instance, while Label dens. reflects the label density, calculated as Label card. divided by the total number of labels.
@@ 6.1.2 Comparison Methods
To evaluate the performance of the proposed method Access-MFS, we compare it with several state-of-the-art multi-label feature selection methods. This includes nine semi-supervised methods (namely SMILE, FSSRDM, SCFS, LEDA, LSMR, SMLFS, SFAM, SMDR-IC, and SFS-BLL), one supervised method (LFFS), and one unsupervised method (UAFS-BH). The following is a brief introduction to the methods being compared.

### v4

@@ 4 Optimization and Algorithm
$\min_{\mathbf{W},\mathbf{F},\mathbf{S},\mathbf{P}}\|\mathbf{H}\mathbf{X}^{T}\mathbf{W}-\mathbf{H}\mathbf{F}\|_{F}^{2}+\lambda\|\mathbf{W}\|_{2,1}+{\|{\mathbf{F}_{l}-\mathbf{Y}_{l}}\|_{F}^{2}}$ (5)
$~~\quad\quad\quad+\theta(\operatorname{Tr}(\mathbf{F}^{T}\mathbf{L}_{s}\mathbf{F})+\operatorname{Tr}(\mathbf{W}^{T}\mathbf{X}\mathbf{L}_{s}\mathbf{X}^{T}\mathbf{W})$
$~~\quad\quad\quad+\alpha\|\mathbf{S}\|_{F}^{2})+\!\mu(\operatorname{Tr}(\mathbf{F}\mathbf{L}_{p}\mathbf{F}^{T})+\beta\|\mathbf{P}\|_{F}^{2})$
@@
$p_{ij}\geq 0,\mathbf{P}\mathbf{1}_{c}=\mathbf{1}_{c}.$
@@ 4.1 Update $\mathbf{W}$ by Fixing Other Variables
Input: Data matrix $\mathbf{X}$, centering matrix $\mathbf{H}$, predicted label matrix $\mathbf{F}$ and Laplacian matrix $\mathbf{L}_{s}$, parameters: $\lambda$ and $\theta$.
@@ 14     Update $\mathbf{W}=\mathbf{R}^{-\frac{1}{2}}\mathbf{A}$;
4.2 Update $\mathbf{F}$ by Fixing Other Variables
$\min_{\mathbf{F}}\|\mathbf{H}\mathbf{X}^{T}\mathbf{W}-\mathbf{H}\mathbf{F}\|_{F}^{2}+\theta\operatorname{Tr}(\mathbf{F}^{T}\mathbf{L}_{s}\mathbf{F})+\mu\operatorname{Tr}(\mathbf{F}\mathbf{L}_{p}\mathbf{F}^{T})$ (12)
$\quad\quad+\|{\mathbf{F}_{l}-\mathbf{Y}_{l}}\|_{F}^{2}.$
$\mathbf{Q}\mathbf{F}+\mu\mathbf{F}\mathbf{L}_{p}=\mathbf{C}$ (14)
@@ 4.3 Update $\mathbf{S}$ by Fixing Other Variables
While $\mathbf{W}$, $\mathbf{F}$ and $\mathbf{P}$ are fixed, the optimization problem in Eq. (4) is equivalent to solve:
$\min_{\mathbf{S}}\operatorname{Tr}(\mathbf{W}^{T}\mathbf{X}\mathbf{L}_{s}\mathbf{X}^{T}\mathbf{W})+\operatorname{Tr}(\mathbf{F}^{T}\mathbf{L}_{s}\mathbf{F})+\alpha\|\mathbf{S}\|_{F}^{2}$ (15)
$\Leftrightarrow$ $\min_{\mathbf{S}}\frac{1}{2}\sum_{i,j=1}^{n}\|\mathbf{W}^{T}\mathbf{x}_{.i}-\mathbf{W}^{T}\mathbf{x}_{.j}\|_{2}^{2}s_{ij}$
$+\frac{1}{2}\sum_{i,j=1}^{n}\|\mathbf{f}_{i.}-\mathbf{f}_{j.}\|_{2}^{2}{s_{ij}}+\alpha\|\mathbf{S}\|_{F}^{2}$
${s.t.}~s_{ii}=0,s_{ij}\geq 0,\mathbf{S}\mathbf{1}_{n}=\mathbf{1}_{n}.$
By setting a matrix $\mathbf{M}=(m_{ij})_{n\times n}$, where $m_{ij}=\frac{1}{2}\|\mathbf{W}^{T}\mathbf{x}_{.i}-\mathbf{W}^{T}\mathbf{x}_{.j}\|_{2}^{2}+\frac{1}{2}\|\mathbf{f}_{i.}-\mathbf{f}_{j.}\|_{2}^{2}$, Eq. (LABEL:New15) is transformed into
@@
In practice, each sample tends to be similar to its neighbours. Therefore, we aim for $\mathbf{s}_{i.}$ to have $k_{s}$ nonzero values, indicating that the $i$th sample is similar to its $k_{s}$ closest neighbours. Assuming that $m_{i1},m_{i2},\cdots,m_{in}$ are arranged in ascending order, we prefer ${s}_{i,k_{s}}>0$ and ${s}_{i,k_{s}+1}=0$. As a result, we have $\alpha=\frac{m_{i,k_{s}+1}}{2\psi}$. Besides, according to the constraint $\mathbf{s}_{i.}\mathbf{1}_{n}=1$, we get $\psi=\frac{1}{k_{s}}+\frac{\sum_{j=1}^{k_{s}}m_{ij}}{2k_{s}\alpha}$. By substituting $\psi$ into the obtained $\alpha$, we can calculate the regularization parameter as $\alpha=\frac{k_{s}m_{i,k_{s}+1}-\sum_{j=1}^{k_{s}}m_{ij}}{2}$. Then, we achieve the optimal solution ${s}_{ij}$ as follows:
${s}_{ij}=\left\{\begin{array}[]{cl}\frac{m_{i,k_{s}+1}+m_{ij}}{k_{s}m_{i,k_{s}+1}-\sum_{h=1}^{k_{s}}m_{ih}}&j\leq k_{s};\\ 0&j>k_{s}.\end{array}\right$. (19)
@@ 4.4 Update $\mathbf{P}$ by Fixing Other Variables
$\Leftrightarrow$ $\min_{\mathbf{P}}\frac{1}{2}\sum_{i,j=1}^{c}\|\mathbf{f}_{.i}-\mathbf{f}_{.j}\|_{2}^{2}p_{ij}+\beta\sum_{i=1}^{c}\|\mathbf{p}_{i.}\|_{2}^{2}$
@@
Following a similar procedure to that used in solving Eq. (17), we can determine the regularization parameter $\beta=\frac{k_{p}g_{i,k_{p}+1}-\sum_{j=1}^{k_{p}}g_{ij}}{2}$, where $k_{p}$ denotes the number of nonzero elements in $\mathbf{p}_{i.}$. Furthermore, the optimal solution ${p}_{ij}$ is formulated as follows:
$p_{ij}=\left\{\begin{array}[]{cl}\frac{g_{i,k_{p}+1}+g_{ij}}{k_{p}g_{i,k+1}-\sum_{h=1}^{k_{p}}g_{ih}}&j\leq k_{p};\\ 0&j>k_{p}.\end{array}\right$. (23)
@@ 5.1 Convergence Analysis
Due to the objective function in Eq. (5) not being convex w.r.t. four variables $\mathbf{W}$, $\mathbf{F}$, $\mathbf{S}$ and $\mathbf{P}$ simultaneously, we address this by dividing it into four sub-objective functions, namely, Eqs. (6), (12), (LABEL:New15), and (20), each of which is separately solved with other variables fixed. Therefore, the convergence of Algorithm 2 can be demonstrated by proving that each sub-objective function decreases monotonically. We first present the following theorem to prove the monotonic decrease of the objective function in Eq. (6) when updating $\mathbf{W}$ with $\mathbf{F}$, $\mathbf{S}$ and $\mathbf{P}$ are fixed.
@@
Proof.
Let $\varUpsilon(\mathbf{W}^{(t)},\mathbf{D}^{(t)})=\|\mathbf{H}\mathbf{X}^{T}\mathbf{W}^{(t)}-\mathbf{H}\mathbf{F}\|_{F}^{2}+\lambda\operatorname{Tr}({\mathbf{W}^{(t)}}^{T}\mathbf{D}^{(t)}\mathbf{W}^{(t)})+\theta\operatorname{Tr}({\mathbf{W}^{(t)}}^{T}\mathbf{X}\mathbf{L}_{s}\mathbf{X}^{T}\mathbf{W}^{(t)})$ represent the objective value in the $t$th iteration of Eq. (7). According to [38], during the alternate updating of $\mathbf{W}$ and $\mathbf{D}$ using Algorithm 1, the following inequalities must hold for the (tt+1)th iteration:
$\varUpsilon(\mathbf{W}^{(t+1)},\mathbf{D}^{(t)})\leq\varUpsilon(\mathbf{W}^{(t)},\mathbf{D}^{(t)})$ (24)
Besides, by means of the property $\|\mathbf{u}\|_{2}-\frac{\|\mathbf{u}\|_{2}^{2}}{2\|\mathbf{v}\|_{2}}\leq\|\mathbf{v}\|_{2}-\frac{\|\mathbf{v}\|_{2}^{2}}{2\|\mathbf{v}\|_{2}}$ for any two nonzero vectors $\mathbf{u}$ and $\mathbf{v}$ as proposed in [37], we have
$\|\mathbf{w}^{(t+1)}_{i.}\|_{2}-\frac{\|\mathbf{w}^{(t+1)}_{i.}\|_{2}^{2}}{2\|\mathbf{w}^{(t)}_{i.}\|_{2}}\leq\|\mathbf{w}^{(t)}_{i.}\|_{2}-\frac{\|\mathbf{w}^{(t)}_{i.}\|_{2}^{2}}{2\|\mathbf{w}^{(t)}_{i.}\|_{2}}$
$\Rightarrow$ $\lambda\sum_{i=1}^{d}\|\mathbf{w}^{(t+1)}_{i.}\|_{2}-\lambda\sum_{i=1}^{d}\frac{\|\mathbf{w}^{(t+1)}_{i.}\|_{2}^{2}}{2\|\mathbf{w}^{(t)}_{i.}\|_{2}}\leq$
$\lambda\sum_{i=1}^{d}\|\mathbf{w}^{(t)}_{i.}\|_{2}-\lambda\sum_{i=1}^{d}\frac{\|\mathbf{w}^{(t)}_{i.}\|_{2}^{2}}{2\|\mathbf{w}^{(t)}_{i.}\|_{2}}$
$\Rightarrow$ $\lambda\|\mathbf{W}^{(t+1)}\|_{2,1}-\lambda\operatorname{Tr}({\mathbf{W}^{(t+1)}}^{T}\mathbf{D}^{(t)}\mathbf{W}^{(t+1)})\leq$
$\lambda\|\mathbf{W}^{(t)}\|_{2,1}-\lambda\operatorname{Tr}({\mathbf{W}^{(t)}}^{T}\mathbf{D}^{(t)}\mathbf{W}^{(t)}).$ (25)
By combing the inequalities (24) and (25), it is easy to obtain
$\|\mathbf{H}\mathbf{X}^{T}\mathbf{W}^{(t+1)}-\mathbf{H}\mathbf{F}\|_{F}^{2}+\lambda\|\mathbf{W}^{(t+1)}\|_{2,1}$ (26)
$+\theta\operatorname{Tr}({\mathbf{W}^{(t+1)}}^{T}\mathbf{X}\mathbf{L}_{s}\mathbf{X}^{T}\mathbf{W}^{(t+1)})$
$\leq\|\mathbf{H}\mathbf{X}^{T}\mathbf{W}^{(t)}-\mathbf{H}\mathbf{F}\|_{F}^{2}+\lambda\|\mathbf{W}^{(t)}\|_{2,1}$
$+\theta\operatorname{Tr}({\mathbf{W}^{(t)}}^{T}\mathbf{X}\mathbf{L}_{s}\mathbf{X}^{T}\mathbf{W}^{(t)}).$
Hence, the objective function in Eq. (5) monotonically decreases by Algorithm 2 in each iteration. ∎
As Eq. (12) is equivalent to Eq. (13), we only need to demonstrate that the objective function in Eq.(13) with respect to $\mathbf{F}$} is a monotonically decreasing function. It is easy to verify that the matrix $\mathbf{Q}$ and the Laplacian matrix $\mathbf{L}_{p}$ are both positive semidefinite. Consequently, the quadratic forms $\mathbf{F}^{T}\mathbf{Q}\mathbf{F}-2\mathbf{F}^{T}\mathbf{C}$ and $\mathbf{F}\mathbf{L}_{p}\mathbf{F}^{T}$ are convex. Hence, the objective function in Eq.(13) is convex. According to [39], we can easily obtain that the objective function in Eq.(13) monotonically decrease by using the updating rules in Algorithm 2. Additionally, the objective functions in Eqs. (LABEL:New15) and (20) are convex w.r.t. $\mathbf{S}$ and $\mathbf{P}$, and possess closed-form solutions, ensuring the convergence of the updates for $\mathbf{S}$ and $\mathbf{P}$, respectively. Therefore, we can conclude that the objective function in Eq.(5) monotonically decreases with each iteration of Algorithm 2. Moreover, the convergence behavior of Algorithm 2 is further demonstrated through experiments.
@@ 5.2 Complexity Analysis
In Algorithm 2, four variables are updated alternately by solving Eqs. (6), (12), (LABEL:New15), and (20). The optimization of Eq. (6) is further transformed into solve Eq. (9). According to [38], the time complexity for updating $\mathbf{W}$ in Eq. (9) is $\mathcal{O}(d^{3})$. For updating $\mathbf{F}$, a proposed algorithm in [39] is applied into solve Eq. (14). Based on [39], we can calculate the computational complexity for updating $\mathbf{F}$, which is $\mathcal{O}(n^{2}c+c^{2}n)$. Eq. (LABEL:New15) is solved by updating $\mathbf{S}$ row by row. According to the closed form solution of $\mathbf{S}$ in Eq. (19), the time complexity of updating $\mathbf{S}$ is $\mathcal{O}(nk_{s}d)$. The optimization $\mathbf{P}$ in Eq. (20) is similar to solving $\mathbf{S}$. In the same way, we can determine the time complexity of $\mathbf{P}$ as $\mathcal{O}(ck_{p}n)$. Therefore, the total computational complexity of Algorithm 2 is $\mathcal{O}((d^{3}+n^{2}c+c^{2}n+nk_{s}d+ck_{p}n)T)$, where $T$ is the number of iteration.
@@ 6.1.1 Datasets
In our experiments, we collected eight real-world benchmark multi-label datasets, each with distinct statistical characteristics, to present the performance of various feature selection methods. These datasets include a biology dataset VirusGO [^1] http://www.uco.es/kdis/mllresources, a medicine dataset Medical1, and six text datasets: Langlog1, Enron1, Slashdot1, Recreation [^2] http://www.lamda.nju.edu.cn/code_MDDM.ashx, Computer2 and, Science2. The detailed statistics for these multi-label datasets are summarized in Table I. Within this table, Label card. refers to the average number of labels per instance, while Label dens. reflects the label density, calculated as Label card. divided by the total number of labels.
@@ 6.1.2 Comparison Methods
To evaluate the performance of the proposed method Access-MFS, we compare it with several state-of-the-art multi-label feature selection methods. These include twelve semi-supervised methods (namely SMILE, FSSRDM, SCFS, LEDA, LSMR, SMLFS, SFAM, SMDR-IC, SFS-BLL, SMDR-MRC, SMFS and SFGR), one supervised method (LFFS), and one unsupervised method (UAFS-BH). The following is a brief introduction to the methods being compared.
@@
$\bullet$ SMDR-MRC [31] incorporates both $\ell_{1}$-norm and $\ell_{2,1}$-norm to extract specific and common features, respectively, and introduces a non-zero correlation constraint term to minimize redundancy and overlapping interference.
$\bullet$ SMFS [32] proposes a novel shared subspace learning framework that decomposes the feature selection matrix into a label-specific weight matrix and a shared subspace projection term, which explicitly captures the correlations among labels.
$\bullet$ SFGR [33] introduces a group decision-making mechanism into feature selection by using the Copeland function to fuse the ranking results of multiple feature subsets, thereby enhancing the stability and robustness of the selection process.